# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

### v1

@@
Abstract.
Traditional domain generalization approaches predominantly focus on leveraging target domain-aware features while overlooking the critical role of source domain-specific characteristics, particularly in federated settings with inherent data isolation. To address this gap, we propose the Federated Source Domain Awareness Framework (FedSDAF), the first method to systematically exploit source domain-aware features for enhanced federated domain generalization (FedDG). The FedSDAF framework consists of two synergistic components: the Domain-Invariant Adapter, which preserves critical domain-invariant features, and the Domain-Aware Adapter, which extracts and integrates source domain-specific knowledge using a Multihead Self-Attention mechanism (MHSA). Furthermore, we introduce a bidirectional knowledge distillation mechanism that fosters knowledge sharing among clients while safeguarding privacy. Our approach represents the first systematic exploitation of source domain-aware features, resulting in significant advancements in model generalization capability.Extensive experiments on four standard benchmarks (OfficeHome, PACS, VLCS, and DomainNet) show that our method consistently surpasses state-of-the-art federated domain generalization approaches, with accuracy gains of 5.2-13.8%. The source code is available at https://github.com/pizzareapers/FedSDAF.
Federated learning, Cross-domain, Federated Domain Generalization, Domain-aware features
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 03–05, 2025; Dublin, Irelandsubmissionid: 5352ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed computing methodologies
Figure 1. Ilustration of Domain Generalization.
[…]
A woman and a girl in white dresses sit in an open car.
1. Introduction
@@
The synergy between federated learning (FL) and large-scale pre-trained models(Houlsby et al., 2019; Su et al., 2024b) has ushered in a new era of distributed learning paradigms for privacy-sensitive applications. While foundational architectures, particularly Transformers, demonstrate remarkable cross-task adaptability(Mahla et al., 2025; Chen et al., 2022) through parameter efficient techniques, conventional FL frameworks(McMahan et al., 2016; Wang et al., 2020) encounter a fundamental limitation: their underlying independent and identically distributed (i.i.d.) assumption stands in stark contrast to the inherent multi-source domain heterogeneity in real-world scenarios(Sahu et al., 2018; Peng et al., 2020; Liu et al., 2021a). This misalignment presents a dual challenge: not only do client-specific data emerge from distinct domain distributions, but models must also simultaneously address inter-client domain shifts while maintaining robust generalization capabilities for unseen target domains during deployment. These intertwined challenges constitute the core problem in Federated Domain Generalization (FedDG).
@@
Standard domain generalization (DG) methods have advanced through Domain-invariant representation learning(Tan et al., 2024; Peng and Chan, 2024) and Feature disentanglement(Cheng et al., 2024). Both Feng et al(Feng et al., 2024a). and Li et al(Li et al., 2024b). have adopted the Feature Disentanglement approach, which separates features into domain-invariant and domain-specific components, utilizing domain-aware features from the target domain to enhance the model’s generalization ability in text classification and target detection tasks. However, these studies primarily focus on leveraging domain-aware features in the target domain, without fully exploring the potential value of domain-aware features in the source domain.
Federated Domain Generalization (FedDG) extends beyond the conventional challenges of standard domain generalization by introducing a critical constraint: the prohibition of raw data sharing across clients. This restriction impedes explicit multi-domain alignment through centralized techniques such as feature matching or adversarial training, thereby limiting the models’ ability to fully leverage knowledge from diverse source domains. Current FedDG approaches have attempted various solutions: Liu et al.(Liu et al., 2021b) facilitated domain knowledge sharing through the exchange of amplitude spectrum information in frequency space; Huang et al.(Huang et al., 2023) implemented class prototype sharing derived from local samples; and Chen et al.(Chen et al., 2023) enabled cross-client style transfer through the exchange of pixel-level feature channel statistics. While these methods represent progress in addressing data isolation, they primarily focus on transferring fragmentary domain characteristics derived directly from local sample statistics, leaving the potential of source domain-aware features largely unexplored.
In the context of federated learning, data isolation fundamentally constrains the model’s ability to effectively utilize knowledge from diverse domains. While existing methods transmit fragmented domain characteristics (e.g., spectral signatures or prototypes), these incomplete knowledge representations fail to capture the crucial relationships between domain-specific patterns and underlying task semantics, resulting in suboptimal cross-domain knowledge integration during model aggregation. Notably, our controlled experiments in Section 3.2 reveal a counterintuitive finding: source domain-aware features demonstrate superior generalization capabilities compared to target domain-derived features in data-isolated scenarios. This observation suggests that the predominant focus of existing methods on target domain adaptation may overlook the valuable contribution of source domain-specific patterns in developing transferable representations.
To address these limitations, we introduce the Federated Source Domain Awareness Framework (FedSDAF), which systematically leverages source domain-aware features through dual adaptive components. The first component, a Domain-Invariant Adapter, preserves essential cross-domain patterns, while the second component, a Domain-Aware Adapter, employs Multi-head Self-Attention (MHSA) to dynamically integrate source-specific knowledge with base model representations, enabling hierarchical knowledge integration. Furthermore, the framework incorporates a bidirectional knowledge distillation mechanism that facilitates logarithmic knowledge exchange between clients and the server, thereby enabling effective knowledge transfer from diverse source domains to the global model. This innovation alleviates the challenges posed by data isolation constraints. Our framework not only enhances the model’s comprehension of domain-specific patterns but also improves cross-domain knowledge integration, ultimately boosting the model’s ability to generalize across multiple target domains. Our contributions are summarized below:
  • We introduce a novel framework, FedSDAF, which systematically addresses the inherent limitations of data isolation in Federated Domain Generalization by leveraging source domain-aware features. The framework consists of dual adaptive components—Domain-Invariant Adapter and Domain-Aware Adapter—that work in tandem to preserve essential cross-domain patterns while dynamically integrating source-specific knowledge. This approach enhances the model’s understanding of domain-specific intricacies, fostering improved generalization capabilities across multiple target domains.
  • We introduce a bidirectional knowledge distillation mechanism that enables knowledge exchange between clients and the server, effectively alleviating the data isolation constraints in federated learning while facilitating comprehensive knowledge transfer from diverse source domains to the global model.Development of a Novel Framework for Leveraging Source Domain-Aware Features.
  • We demonstrate the framework’s effectiveness through extensive experiments on four benchmark datasets, achieving state-of-the-art performance in domain generalization tasks.
2. Related Work
2.1. Federated Learning
Federated Learning represents a paradigm shift in distributed model optimization, enabling collaborative training across multiple devices or organizations without sharing raw data(Tang and Wang, 2023; Fu et al., 2024). The fundamental principle involves local model training on client devices followed by parameter aggregation at a central server, creating a global model that benefits from diverse data sources while maintaining data privacy(Gad et al., 2024; Cai et al., 2024). One of the primary challenges in federated learning environments is the non-Independent and Identically Distributed (non-IID) nature of client data, which significantly impacts model performance(Islam et al., 2024; Feng et al., 2024b). To address these challenges, researchers have proposed various schemes and models. For instance, Zhang et al(Zhang et al., 2022). introduced a method for personalizing local models, while Guo et al(Guo and Yang, 2024). developed a more robust algorithm aimed at learning more efficient global models(Li et al., 2024a). Despite these advances, most research on non-IID challenges focuses on addressing out-of-sample gaps in unseen client data with known distributions. However, as Yuan et al. (Yuan et al., 2021) highlight, there is still a critical gap in handling unseen client distributions(Jiang et al., 2024). This highlights the need for approaches that can generalize beyond the training distributions, especially when encountering entirely new domains.
2.2. Domain Generalization
The primary objective of domain generalization is to enable models to perform well on unseen domains by learning robust representations from multiple source domains. Within this framework, representation learning has emerged as a pivotal approach, concentrating on the extraction of domain-agnostic features that generalize across different distributions(Lv et al., 2022; Liu et al., 2024a). Existing research in representation learning for domain generalization can be broadly categorized into two main strands: Domain-Invariant Representation Learning (DIRL)(Xu et al., 2022) and Feature Disentanglement(Gholami et al., 2023). DIRL methods aim to extract features that remain consistent across domains through techniques such as kernel tricks, adversarial learning, and explicit feature alignment. These approaches strive to minimize the distribution discrepancy between source domains, operating under the assumption that domain-invariant representation will transfer effectively to unseen target domains(Su et al., 2024a). In contrast, feature disentanglement involves dividing input features into domain-invariant and domain-specific components. This separation allows models to isolate generalizable patterns from domain-specific variations, potentially enhancing performance on unseen domains. Significant contributions in this area include the work of Feng et al(Feng et al., 2024a). and Li et al(Li et al., 2024b)., who employed feature disentanglement to improve model generalization in text classification and target detection tasks. Additionally, they used features specific to the target domain, combined with domain-invariant features, to improve the model’s performance on that domain. However, a notable limitation of current feature disentanglement approaches is their predominant focus on utilizing domain-aware features from the target domain, while not fully exploring the potential value of domain-aware features from source domains(Zhang et al., 2023a). This represents a significant opportunity, particularly in federated scenarios where data isolation presents additional challenges.
2.3. Federated Domain Generalization
Federated Domain Generalization (FedDG) is an emerging research area that addresses the challenges posed by traditional federated learning (FL) and domain generalization (DG) approaches(Zhang et al., 2023d). The primary objective of FedDG is to tackle data isolation in federated scenarios while enhancing the generalization capabilities of models across unseen target domains(Radwan and Shehata, 2024). For instance, Liu et al(Liu et al., 2021b). proposed a method that leverages the exchange of frequency-space magnitude spectral information. Similarly, Wan et al(Wan et al., 2024). facilitated the sharing of class prototypes extracted from local samples among different clients. Furthermore, Chen et al. enabled cross-client style migration by exchanging the mean and variance of pixel-level feature channels. Recent studies, such as FedGCA(Liu et al., 2024b), have also introduced global consistency augmentation to address the challenges associated with single-source federation domain generalization. Although these approaches mitigate the issue of data segregation to some extent, they still exhibit significant limitations, The exchanged domain knowledge components (spectral signatures, prototypes, style parameters) operate at shallow feature levels, failing to capture deep semantic correlations between domain-specific patterns and task objectives.
3. Methodology
3.1. Problem Setup
Traditional Domain Generalization (DG): Consider $K$ source domains ${\mathcal{S}^{k}}{k=1}^{K}$ with distinct distributions $P^{k}\neq P^{k^{\prime}}\ (k\neq k^{\prime})$, where each domain consists of labeled samples $\mathcal{S}^{k}={(x_{i}^{k},y_{i}^{k})}{i=1}^{n_{k}}$. The objective is to learn a model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ that minimizes the expected risk on an unseen target domain $\mathcal{T}$ with distribution $P^{\text{tgt}}$:
(1) $\min_{\theta}\mathbb{E}{(x,y)\sim P^{\text{tgt}}}\mathcal{L}(f\theta(x),y)$
Federated Domain Generalization (FedDG): Extends the traditional DG framework by incorporating a critical privacy constraint :
Data Isolation: Source domains are distributed across $K$ clients ${\mathcal{C}k}{k=1}^{K}$, where each client’s data must remain private and locally contained.
The FedDG objective can be formulated as a constrained optimization problem:
(2) $\begin{split}\min_{\theta_{g}}&\underbrace{\mathbb{E}_{(x,y)\sim P^{\text{tgt}}}\mathcal{L}(f_{\theta_{g}}(x),y)}_{\text{Target Risk}}\\ &\text{s.t.}\quad\underbrace{\theta_{g}=\mathcal{A}(\theta_{k}),\ \mathcal{S}^{k}\subseteq\mathcal{C}_{k}}_{\text{Federated Constraints}}\end{split}$
3.2. Motivational study
Building upon the problem setup of federated domain generalization , we pose a critical research question: Which type of domain-aware features contributes more significantly to model performance in target domains - those derived from the source domain or the target domain itself? To investigate this, we conduct controlled experiments on the PACS dataset using a leave-one-domain-out protocol. Specifically, we train domain-aware adapters independently on three source domains (Photo, Art, Cartoon) while keeping the domain-invariant adapter fixed. Each adapter is then evaluated on all possible target domains without any target domain fine-tuning. Table 1 reveals a counterintuitive finding: when testing on a target domain (e.g., Sketch), adapters trained on other source domains consistently outperform those trained on the target domain itself. For instance, the Art-trained adapter achieves 97.07% accuracy on Sketch, surpassing the Sketch-trained adapter’s 94.94%.
Target Domain P A C
Adapter in Domain P 96.49% 97.07% 97.82%
Adapter in Domain A 98.86% 89.42% 97.35%
Adapter in Damain C 98.26% 94.38% 94.94%
Table 1. Accuracy results of the domain-aware perceptual adapters combined with the global domain-invariant adapter across different target domains in the PACS dataset.
This phenomenon can be explained through two key perspectives from our introduction: (1) Data isolation constraints in federated learning prevent effective collaborative modeling of target domain features, making their isolated extraction suboptimal; (2) Source domain features provide complementary contextual knowledge that helps bridge domain gaps through their inherent diversity. These findings motivate our core hypothesis: Systematic exploitation of source domain-aware features through collaborative federation can create more robust representations for unseen target domains. This insight drives our framework design that strategically aggregates source domain knowledge while preserving domain-specific characteristics through novel adaptation mechanisms.
3.3. Proposed Method
@@
Figure 2. FedSDAF Architecture Design.
[…]
A woman and a girl in white dresses sit in an open car.
3.3.1. Knowledge Integrator Adapter (KIA)
@@
(3) $\hat{A}_{dw}\left(h\right)=MHSA\left(LN\left[A_{dw}\left(h\right);h\right]\right)$
(4) $h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$
@@
Figure 3. Optimization process
3.3.2. Bidirectional Knowledge Distillation (BKD)

### v3

@@
Abstract
Traditional Federated Domain Generalization (FedDG) methods focus on learning domain-invariant features or adapting to unseen target domains, often overlooking the unique knowledge embedded within the source domain, especially in strictly isolated federated learning environments. Through experimentation, we discovered a counterintuitive phenomenon.: features learned from a complete source domain have superior generalization capabilities compared to those learned directly from the target domain. This insight leads us to propose the Federated Source Domain Awareness Framework (FedSDAF), the first systematic approach to enhance FedDG by leveraging source domain-aware features. FedSDAF employs a dual-adapter architecture that decouples "local expertise" from "global generalization consensus". A Domain-Aware Adapter, retained locally, extracts and protects the unique discriminative knowledge of each source domain, while a Domain-Invariant Adapter, shared across clients, builds a robust global consensus. To enable knowledge exchange, we introduce a Bidirectional Knowledge Distillation mechanism that facilitates efficient dialogue between the adapters. Extensive experiments on four benchmark datasets (OfficeHome, PACS, VLCS, DomainNet) show that FedSDAF significantly outperforms existing FedDG methods.The source code is available at https://github.com/pizzareapers/FedSDAF.
1 Introduction
@@
The synergy between federated learning (FL) and large-scale pre-trained models [2, 3] has ushered in a new era of distributed learning paradigms for privacy-sensitive applications. While foundational architectures, particularly Transformers, demonstrate remarkable cross-task adaptability [1, 4] through parameter efficient techniques, conventional FL frameworks [5, 6] encounter a fundamental limitation: their underlying independent and identically distributed (i.i.d.) assumption stands in stark contrast to the inherent multi-source domain heterogeneity in real-world scenarios [7, 8, 9]. This misalignment presents a dual challenge: not only do client-specific data emerge from distinct domain distributions, but models must also simultaneously address inter-client domain shifts while maintaining robust generalization capabilities for unseen target domains during deployment. These intertwined challenges constitute the core problem in Federated Domain Generalization (FedDG).
@@
Standard domain generalization (DG) methods have advanced through Domain-invariant representation learning [14, 15] and Feature disentanglement [16]. Both Feng et al. [12] and Li et al. [13] have adopted the Feature Disentanglement approach, which separates features into domain-invariant and domain-specific components, utilizing domain-aware features from the target domain to enhance the model’s generalization ability in text classification and target detection tasks. However, these studies primarily focus on leveraging domain-aware features in the target domain, without fully exploring the potential value of domain-aware features in the source domain.
The core challenge in Federated Domain Generalization (FedDG) is the strict data isolation among clients, which impedes the model’s ability to align knowledge from diverse source domains [43]. While existing methods attempt to circumvent this by exchanging fragmentary information like spectral data, class prototypes, or style statistics [43, 10, 11], these incomplete representations of knowledge lead to suboptimal fusion during model aggregation.
However, our motivational study reveals a counter-intuitive finding that challenges the conventional focus on domain invariance: in isolated settings, source domain-aware features possess superior generalization capabilities. For instance, on the PACS dataset, an adapter trained solely on the "art painting" source domain achieved 97.07% accuracy on the unseen "sketch" target, significantly outperforming an adapter trained and tested on "sketch" itself (94.94%) [21, 23]. The underlying logic is that a source-aware model, forced to bridge a significant domain gap, must discard superficial source styles (e.g., brushstrokes) and instead learn abstract, essential semantic features, which inherently generalize better to any unseen domain.
This insight reframes the problem: the true challenge is not to eliminate domain differences but to systematically leverage the unique knowledge from each source domain. This creates a core contradiction between preserving client-specific "local expertise" and building a unified "global consensus," a task where simplistic aggregation methods like FedAvg inevitably fail by diluting critical local knowledge.
To resolve this tension, we propose the Federated Source Domain-Aware Framework (FedSDAF). The innovative core of FedSDAF is a unique dual-adapter architecture that transforms this contradiction into a symbiotic relationship through principled decoupling and collaboration. We introduce a fully private, locally-retained Domain-Aware Adapter to act as a "local expert," tasked with mining and safeguarding unique source domain knowledge. In parallel, a shared Domain-Invariant Adapter serves as a "global generalizer," building a robust consensus across all clients. These two roles are interconnected by a novel Bidirectional Knowledge Distillation (BKD) mechanism. This BKD process enables a synergistic dialogue: the local adapter enhances the global model’s generalization via its distilled knowledge, while the global model in turn guides the local adapter to prevent overfitting. It is this dynamic interplay between local expertise and global consensus that allows FedSDAF to systematically transform source-specific knowledge from overlooked "noise" into a core "signal" that fundamentally enhances generalization under data isolation.
In this way, FedSDAF transforms source domain knowledge from the "noise" traditionally overlooked by existing methods into a core "signal" that enhances generalization capabilities, systematically addressing the challenges posed by data isolation. Our contributions can be summarized as follows:
  • Proposing a novel FedDG paradigm: We reveal and validate through rigorous experiments that, in a federated environment with data isolation, source domain-aware features possess greater generalization potential than target domain features. Based on this core insight, we introduce FedSDAF, which systematically utilizes source domain knowledge to enhance model generalization by principled decoupling and collaboration of local domain expertise and global generalization consensus.
  • Designing an innovative collaborative learning architecture: Our dual-adapter and bidirectional knowledge distillation mechanism constructs a dynamic collaborative system. In this system, a private "local expert" adapter and a shared "global generalizer" adapter can mutually promote and co-evolve through a knowledge dialogue mechanism, effectively integrating rich local insights into the global model.
  • Achieving state-of-the-art performance on multiple benchmarks: We conducted extensive experiments on four challenging domain generalization benchmark datasets. The results demonstrate that by systematically harnessing source domain awareness, FedSDAF significantly outperforms existing methods, achieving new state-of-the-art performance and robustly validating the effectiveness and superiority of our proposed approach.
2 Related Work
2.1 Federated Learning and Domain Generalization
Federated Learning (FL) enables collaborative model training on decentralized data without sharing raw information [40, 55, 35, 56]. A primary challenge in FL is the non-Independent and Identically Distributed (non-IID) nature of client data, which impairs model performance [49, 57]. While existing methods address this via personalization [51] or robust aggregation [52, 58], they often struggle to generalize to entirely unseen distributions [19, 59], a core problem addressed by Domain Generalization (DG).
Standard DG learns robust representations from multiple source domains, often via Domain-Invariant Representation Learning (DIRL) [46, 38, 60, 37] or Feature Disentanglement [47]. These approaches, however, predominantly focus on leveraging features from the target domain [12, 13], leaving the potential of source-domain-aware features largely unexplored, especially in federated settings [48].
2.2 Federated Domain Generalization
Federated Domain Generalization (FedDG) aims to enhance model generalization to unseen target domains under the data isolation constraints of FL [41, 42]. Prior FedDG studies have explored sharing frequency-space information [43], class prototypes [44], pixel-level channel statistics [11], or using global consistency augmentation [45]. While useful, these methods often exchange shallow feature components (e.g., spectral signatures, style parameters), failing to capture the deeper semantic correlations necessary for robust generalization. Our work addresses this gap by systematically exploiting source domain-aware features.
3 Methodology
3.1 Problem Setup
In the Federated Domain Generalization (FedDG) setting, we consider a scenario involving $K$ independent clients. Each client, $C_{k}$, exclusively holds a local source dataset $D_{k}=\{(x_{ik},y_{ik})\}_{i=1}^{n_{k}}$ that is drawn from a unique probability distribution $P_{k}(X,Y)$, such that $P_{k}\neq P_{j}$ for any $k\neq j$. The foundational constraint of this setting is: Data Isolation, raw data sharing among clients is strictly prohibited due to privacy concerns. Clients can only collaborate to train a model by exchanging non-sensitive information, such as model parameters or updates, via a central server. The primary objective in FedDG is to leverage the diverse, isolated source domains $\{D_{k}\}_{k=1}^{K}$ to learn a single, robust global model $f_{\theta_{g}}$ that performs well on an entirely unseen target domain $D_{t}$. This target domain is drawn from a different distribution $P_{t}$, where $P_{t}\neq P_{k}$ for all source clients $k\in\{1,\ldots,K\}$. The global model $f_{\theta_{g}}$ is produced by aggregating the local models trained on each client’s private data. This constrained optimization problem is formally expressed as minimizing the expected risk on the target domain:
$\theta_{g}^{*}=\min_{\theta_{g}}\mathbb{E}_{(x,y)\sim P_{t}}\left[L(f_{\theta_{g}}(x),y)\right]$
where $L$ represents a predefined loss function.
3.2 Motivational Study: The Case for Source-Domain Awareness
The central hypothesis of our work is that in data-isolated settings, features learned from a complete source domain offer a more generalizable foundation than features learned from the target domain itself. To validate this, we conduct a carefully designed motivational study.
Experimental Setup. We employ a strict leave-one-domain-out protocol on the PACS dataset. For each pair of a single source domain $\mathcal{D}_{S}$ and a single target domain $\mathcal{D}_{T}$ (where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$), we conduct two parallel and fair experiments:
  1. 1.
    Source-Aware Generalization: An adapter is trained only on the training set of the source domain $\mathcal{D}_{S}$ (e.g., Art Painting) and is evaluated on the test set of the unseen target domain $\mathcal{D}_{T}$ (e.g., Sketch). This measures the ability to generalize knowledge across a significant domain shift.
  2. 2.
    Target-Aware In-Domain Generalization: A separate adapter is trained only on the training set of the target domain $\mathcal{D}_{T}$ (e.g., Sketch) and is evaluated on its own test set. This represents the standard in-domain performance baseline.
This one-to-one setup ensures that both adapters are trained on datasets of comparable size and diversity, isolating the effect of the learning paradigm itself.
Table 1: Performance (%) of adapters trained on a single source domain (rows) and evaluated on a single target domain (columns).
Adapter Trained on Target Domain (Y)
(X) Photo Art Cartoon Sketch
Photo 96.55 98.92 98.14 94.21
Art Painting 98.88 98.11 98.25 97.07
Cartoon 98.03 98.67 97.23 95.33
Sketch 97.51 97.93 94.98 94.94
Results and Principled Interpretation. The results, presented in Table 1, reveal a clear and theoretically grounded trend. In almost all cases, the Source-Aware adapter (off-diagonal values) outperforms the target-aware adapter (diagonal values) when evaluated on the target domain. For instance, when targeting the Sketch domain, the adapter trained on Art Painting (97.07%) significantly surpasses the one trained on Sketch itself (94.94%).
This seemingly counter-intuitive outcome stems from a fundamental learning principle. The Target-Aware adapter is prone to learning spurious correlations by associating domain-specific styles (e.g., pencil texture in Sketch) with class labels, leading to a brittle representation that overfits. In contrast, the Source-Aware adapter, forced to generalize across a wide domain gap, cannot rely on the source style (e.g., brushstrokes). This constraint compels the optimization to discard superficial features and instead learn more abstract, semantically invariant representations that capture the essence of the classes. This robust semantic foundation, stripped of domain-specific biases, naturally exhibits superior generalization.
3.3 Proposed Method
@@
Figure 1: FedSDAF Architecture Design.
3.3.1 Knowledge Integrator Adapter (KIA)
@@
$\hat{A}_{dw}\left(h\right)=MHSA\left(LN\left[A_{dw}\left(h\right);h\right]\right)$ (1)
$h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$ (2)
@@
Figure 2: Optimization process.
3.3.2 Bidirectional Knowledge Distillation (BKD)